# Leveraging Convolutional and Graph Networks for an Unsupervised Remote Sensing Labelling Tool

### v1

@@ Abstract
Machine learning for remote sensing imaging relies on up-to-date and accurate labels for model training and testing. Labelling remote sensing imagery is time and cost intensive, requiring expert analysis. Previous labelling tools rely on pre-labelled data for training in order to label new unseen data. In this work, we define an unsupervised pipeline for finding and labelling geographical areas of similar context and content within Sentinel-2 satellite imagery. Our approach removes limitations of previous methods by utilising segmentation with convolutional and graph neural networks to encode a more robust feature space for image comparison. Unlike previous approaches we segment the image into homogeneous regions of pixels that are grouped based on colour and spatial similarity. Graph neural networks are used to aggregate information about the surrounding segments enabling the feature representation to encode the local neighbourhood whilst preserving its own local information. This reduces outliers in the labelling tool, allows users to label at a granular level, and allows a rotationally invariant semantic relationship at the image level to be formed within the encoding space.
@@ 1 Introduction
Rapid advancements in satellite imagery and Remote Sensing (RS) have led to an explosion of data, with the Sentinel-2 mission producing 1.6 terabytes of data daily. This wealth of data enables applications across agriculture, land use mapping, environmental monitoring, coastal management and monitoring and geological studies (Karthikeyan et al., 2020; Zhang and Zhang, 2022; Weiss et al., 2020; Ahmadi and Pekkan, 2021; Seale et al., 2022) – all of which require high quality labelled datasets. The relatively large pixel resolutions of Sentinal-2 (covering 10 to 60m) create challenges for labelling. A single pixel, representing a 10m by 10m area, might cover both land and water, leading to significant variation in the quantities of each within the same pixel. Dataset labels can consist of small image segments that have a feature of interest present, labelled bounding boxes or pixel level labels.
There is a pressing need for accurate labelled datasets in remote sensing to enable reliable analysis and model training. A notable gap in remote sensing tools for labelling was identified by Li et al. (Li et al., 2021a), who developed LabelRS – a segmentation and dataset-building tool. LabelRS leverages open-source data from OpenStreetMaps, Land Use Land Cover (LULC) datasets, and ArcGIS user annotations. While it is effective for generating semantic, object, or image classification datasets, its reliance on predefined label categories limits flexibility, making it less adaptable to new or unseen categories.
To address this challenge, we sought to develop a toolchain that enhances labelling efficiency by presenting experts with similar content for quick and accurate annotation. We turned to the related field of Content-Based Image Retrieval (CBIR), which aims to retrieve similar content through three main steps: 1) feature representation, 2) indexing, and 3) similarity measurement (Li et al., 2021b). By leveraging CBIR’s ability to present relevant images in an unsupervised context, we can offer increased flexibility and time efficiency, assuming the retrieved content is sufficiently similar. Several image retrieval tools have adopted CBIR to return a list of images most similar to a target query, such as those described in (Qi et al., 2020; Hou et al., 2019).
@@
CBIR methods typically begin with low-level feature extractors that focus on local texture, colour (spectral), and shape characteristics. Examples of such feature extraction techniques include the Scale-Invariant Feature Transform (SIFT) (Chen et al., 2020; Zhu et al., 2016), Gray-Level Co-occurrence Matrices (GLCM), wavelets, Gabor filters, and Local Binary Patterns (LBP) (Huang et al., 2014; Ranchin and Wald, 1993; Yang and Newsam, 2008; Huang et al., 2016). These techniques primarily focus on either texture features (which capture spectral relationships) or the spectra themselves. Once extracted, these low-level features can be further processed into mid-level representations using algorithms like the Vector of Locally Aggregated Descriptors (VLAD), Bag of Visual Words (BoW), or Bag of Features (Jégou et al., 2010; Sivic and Zisserman, 2003; Cao et al., 2010). BoW is a foundational technique that employs K-means clustering to create a visual codebook, which then generates histograms of local features. While widely used, BoW is often outperformed by VLAD, which improves upon it by capturing the spatial relationships between local features and their corresponding cluster centres.
Deep Convolutional Neural Networks (DCNNs) (Song et al., 2019) can replace traditional low-level feature extractors, leading to substantial improvements in image retrieval performance (Imbriaco et al., 2019; Bosilj et al., 2016). For instance, Wang et al. (Wang et al., 2017) employ multiple convolutional layers to build hierarchical feature sets, which are then enhanced using VLAD descriptors. Their work demonstrates the efficacy of DCNN architectures such as VGG, ResNet, OverFeat, and CaffeNet in this context. More advanced models like NetVLAD and SatResNet-50 outperform traditional hand-crafted descriptors in terms of retrieval accuracy (Napoletano, 2018). Additionally, these DCNN-based methods can function in an unsupervised manner, utilising techniques like autoencoders, generative models, and a combination of support vector machines (SVMs) or visual codebooks (Romero et al., 2015; Tang et al., 2018; Sun and Bourennane, 2020). Wang et al. (Wang et al., 2021) propose a deep hashing approach using Graph Neural Networks (GNNs), where deep features are extracted from a pre-trained ResNet network.
Despite their effectiveness, a limitation of visual codebook-based methods is that they often lose semantic relationships between image features. To overcome this, Sumbul and Demir (Sumbul and Demir, 2021) explore multi-label co-occurrence relationships in remote sensing (RS) images. They employ a CNN feature extractor, which feeds into a Graph Neural Network (GNN) that compares features using multiple Siamese Neural Networks (SNN), each with a different form of triplet loss to refine the results. Yu et al. (Yu et al., 2022) leverage GNNs for cross-model text-image matching, showing significant improvements when combined with attention mechanisms. Similarly, Kang et al. (Kang et al., 2021) enhance multi-label image retrieval by using graph structures to better represent the relationships between multiple semantic annotations.
By better organising the feature space, these methods can improve matching accuracy. Explicit feature space manipulation techniques, such as triplet loss (Zhang et al., 2021; Yan et al., 2020; Draganov and Dohn, 2023), have shown promise in this area. Halladin-Dąbrowska et al. (Halladin-Dąbrowska et al., 2019) used t-SNE to enhance reference data for heterogeneous non-forest vegetation. Their approach, which employs unsupervised dimensionality reduction, differs from previous methods by eliminating the need for predefined label categories – an issue in many remote sensing applications where label definitions vary widely. Additionally, their method provides contextual feedback about the relationships between features, such as distinguishing urban areas from surrounding land or dense urban zones. This contextual discrimination, based on geographical proximity, allows labellers to customise datasets more effectively, ensuring they can balance and fine-tune their labelled datasets according to specific requirements.
@@
Our research addresses the gap between low-level features and high-level semantic concepts, as identified by Chaudhuri et al. (Chaudhuri et al., 2018). They proposed a sub-graph matching strategy that emphasizes neighbourhood contexts within segmented images. While effective, this approach is computationally expensive (Chaudhuri et al., 2018). More recent work (Chaudhuri et al., 2022) improves on this by using contrastive loss between Siamese Graph Convolutional Networks (GCNs), replacing sub-graph matching. This method applies context-based attention to nodes, enhancing image retrieval performance. The key advantage of GCNs is their ability to encode neighbourhood information into each segment through message-passing operations, facilitating more context-aware feature extraction.
Patel et al. (Patel et al., 2023) propose a two-dimensional embedding approach to facilitate mass labelling. Their method first generates features using an Autoencoder (AE), followed by dimensionality reduction techniques such as t-SNE and UMAP. These techniques are useful for preserving local distances between data points when reducing dimensionality, which aids in visualising machine learning results (Ali et al., 2019, 2021). In particular, they employ these embeddings to extract, visualise, and label features from images in a two-dimensional space, allowing for easy exploration via interactive tools like pan, zoom, and brushing. However, their approach struggles with certain outliers. For instance, in their study, the orientation and intensity of textural features caused clouds in satellite imagery to be incorrectly clustered with land and water borders, leading to inaccuracies in labelling.
Patel et al.’s method is limited by issues of feature misalignment due to a lack of contextual understanding. In contrast, we aim to improve the robustness of feature embeddings by integrating Graph Neural Networks (GNNs) and context-based attention mechanisms, which encode neighbourhood relationships and enable the creation of more accurate, context-aware embeddings for satellite imagery. Our method mitigates the impact of outliers and provides a more reliable tool for visual labelling, ultimately enhancing both the quality and precision of dataset creation.
In this paper, we propose an enhanced unsupervised feature extraction method for visual labelling and exploration. Specifically, we make the following contributions:
@@
    By integrating Graph Neural Networks (GNNs) and feature encoding on image segmentations, we demonstrate that the resulting two-dimensional embedding space is rotationally invariant, making the approach more robust to spatial transformations.
@@
    Finally, we apply this methodology to develop an advanced interactive tool for exploration and labelling, enhancing both efficiency and accuracy in labelling tasks.
@@ 2 Materials and Methods
The proposed pipeline consists of three key components (Figure 1): a U-Net for feature extraction, a Graph Neural Network (GNN) for generating relational embeddings, and a graph-matching algorithm for assessing graph similarity. The U-Net, like most convolutional networks, learns a generalised representation of the input data. The goal is to enable the U-Net to process unseen images and produce activation maps that generalise well for downstream tasks. The next component in the pipeline is the GNN, which incorporates geographical context by capturing neighbourhood information. For example, it distinguishes between a water pixel surrounded by ocean versus one near a beach coastline. To convert the U-Net’s activation maps into a graph structure suitable for the GNN, segmentation (into super-pixels) is applied. Finally, graph-matching algorithms are used for dimensionality reduction and to quantify the similarity between graphs.
@@ 2.1 Dataset
We utilise the SWED dataset (Seale et al., 2022), compiled of Sentinel-2 images captured 2017–2021. The images are tiles, regions of 10,980m by 10,980m, taken in a single pass of the satellite. Minimal cloud coverage is present within the dataset, and each tile contains a diverse set of coastal landscape features from globally distributed locations.
@@ 2.2 Fuzzy C-Means target extraction
Due to the spatial resolution of most satellite platforms, pixel assignment can be ambiguous or uncertain (Zhang et al., 2024). Fuzzy C-means extraction, unlike its K-means counterpart, assigns to each data point a degree of membership to each cluster. This likelihood per pixel for each class transforms the problem domain into a multi-label classification. We use all 12 channels for the training set and to reduce memory, we use subsampling during model training, where we sample every $56^{th}$ pixel. Finally, we applied the clustering model to the entire image to attain a tile-wide classification. The resultant centroids do not necessarily relate to certain physical materials such as water or concrete, more so a representative of the most common pixel properties. The soft classification allows for generalisation of data that does not attribute fully to any centroid.
@@ 2.3 Pre-Proccessing
For processing by the U-Net and GNN we split the tile into chips. A chip is created by applying a sliding window of size $\times 256$ 256 pixels to the larger image, with no overlap between consecutive windows. This results in $\times 42$ 42 chips containing $752\times 10,$ 752 × 10 , 752 of the original tiles $980\times 10,$ 980 × 10 , 980 pixels. For training and testing, we split the dataset by taking every fourth chip in a top-down, row-by-row pattern for the test set obtaining a 75/25 split. The size of the chip is a balance between training size and preserving contextual neighbourhood information. The larger the image the more computationally expensive the resulting pipeline becomes however more information about neighbouring features can be encoded. Considering CNN filters, larger chips result in fewer edge cases.
@@ 2.4 CNN feature extraction
To learn the texture representation we utilise a CNN where the learning function considers the C-Means classification for an objective discrimination of features. The CNN of choice was a U-Net, with 5 layers each for the encoder and decoder. Starting with 64 kernels in the encoder, these are doubled for each consecutive layer to 1024 in the final layer and opposite for the decoder. The target of the model is the fuzzy C-Means predictions. We used 2, 8 and 18 clusters to explore the performance of the U-Net. As each fuzzy classification is non-exclusive, a multi-label loss function is used. The loss function combined Dice loss and binary cross entropy Seale et al. (2022).

### v2

@@ Abstract
Machine learning for remote sensing imaging relies on up-to-date and accurate labels for model training and testing. Labelling remote sensing imagery is time and cost intensive, requiring expert analysis. Previous labelling tools rely on pre-labelled data for training in order to label new unseen data. In this work, we define an unsupervised pipeline for finding and labelling geographical areas of similar context and content within Sentinel-2 satellite imagery. Our approach removes limitations of previous methods by utilising segmentation with convolutional and graph neural networks to encode a more robust feature space for image comparison. Unlike previous approaches we segment the image into homogeneous regions of pixels that are grouped based on colour and spatial similarity. Graph neural networks are used to aggregate information about the surrounding segments enabling the feature representation to encode the local neighbourhood whilst preserving its own local information. This reduces outliers in the labelling tool, allows users to label at a granular level, and allows a rotationally invariant semantic relationship at the image level to be formed within the encoding space. Our pipeline achieves high contextual consistency, with similarity scores of SSIM = 0.96 and SAM = 0.21 under context-aware evaluation, demonstrating robust organisation of the feature space for interactive labelling.
@@ 1 Introduction
Rapid advancements in satellite imagery and Remote Sensing (RS) have led to an explosion of data, with the Sentinel-2 mission producing 1.6 terabytes of data daily. This wealth of data enables applications across agriculture, land use mapping, environmental monitoring, coastal management and monitoring and geological studies (Karthikeyan et al., 2020; Zhang and Zhang, 2022; Weiss et al., 2020; Ahmadi and Pekkan, 2021; Seale et al., 2022) – all of which require high quality labelled datasets. However, the relatively large pixel resolutions of Sentinal-2 (covering 10 to 60m) create challenges for labelling. A single pixel, representing a 10m by 10m area, may cover both land and water, introducing ambiguity and mixed spectral responses. Traditional labelling approaches, whether bounding boxes, pixel-level masks, or small image segments, are time consuming and costly, requiring expert input.
Existing Labelling Tools
Several tools have been developed to address these challenges. For example, Li et al. introduced LabelRS (Li et al., 2021a), which leverages OpenStreetMaps, Land Use Land Cover (LULC) datasets, and ArcGIS annotations to generate semantic, object, or image classification datasets. While effective, such tools rely on predefined label categories, limiting flexibility and adaptability to new or unseen classes. Beyond LabelRS, most existing solutions are not full labelling platforms but retrieval-based systems that assist annotation by retrieving and presenting similar images to a target image. These include CBIR-driven frameworks (Qi et al., 2020; Hou et al., 2019) and deep learning extensions such as NetVLAD and SatResNet-50 (Napoletano, 2018), which improve retrieval accuracy but still depend on prior labels for supervised training. More recently, Patel et al. (Patel et al., 2023) proposed an interactive labelling tool based on manifold embeddings (UMAP and t-SNE), enabling experts to explore and annotate clusters in a 2D space. However, this approach struggles with outliers and lacks contextual awareness, highlighting the need for more robust, context-sensitive labelling solutions.
Methods Used in These Tools
To improve labelling efficiency, many approaches borrow from Content-Based Image Retrieval (CBIR), which retrieves similar content through three main steps: feature representation, indexing, and similarity measurement (Li et al., 2021b). CBIR-based systems aim to present visually similar images to assist annotation, assuming that retrieved content shares sufficient semantic similarity.
Traditional Feature Extraction.
@@
Early CBIR methods relied on low-level descriptors capturing texture, colour, and shape characteristics. Common techniques include Scale-Invariant Feature Transform (SIFT) (Chen et al., 2020; Zhu et al., 2016), Gray-Level Co-occurrence Matrices (GLCM), wavelets, Gabor filters, and Local Binary Patterns (LBP) (Huang et al., 2014; Ranchin and Wald, 1993; Yang and Newsam, 2008; Huang et al., 2016). These techniques primarily focus on either texture features (which capture spectral relationships) or the spectra themselves. Once extracted, these descriptors are often aggregated into mid-level representations using algorithms such as Bag of Visual Words (BoW) and Vector of Locally Aggregated Descriptors (VLAD) (Jégou et al., 2010; Sivic and Zisserman, 2003; Cao et al., 2010). BoW employs K-means clustering to create a visual codebook and histograms of local features, while VLAD improves upon BoW by encoding spatial relationships between features and cluster centres. Despite their popularity, these methods typically lose semantic relationships between image features, limiting their effectiveness for complex remote sensing imagery.
Deep Learning-Based Retrieval.
Deep Convolutional Neural Networks (DCNNs) have transformed image retrieval by learning hierarchical features that outperform handcrafted descriptors (Song et al., 2019; Imbriaco et al., 2019). Architectures such as VGG, ResNet, OverFeat, and CaffeNet have been widely adopted (Wang et al., 2017), with extensions like NetVLAD and SatResNet-50 achieving state-of-the-art performance in remote sensing retrieval tasks (Napoletano, 2018). These models aggregate convolutional features into compact descriptors, often combined with VLAD encoding for improved spatial awareness (Wang et al., 2017). Unsupervised variants using autoencoders, generative models, and hybrid approaches with SVMs or visual codebooks further reduce reliance on labelled data (Romero et al., 2015; Tang et al., 2018; Sun and Bourennane, 2020). However, even deep models can struggle with contextual relationships when features are treated independently.
Graph-Based and Context-Aware Methods.
@@
To address the semantic gap, recent work incorporates Graph Neural Networks (GNNs) to encode spatial context and multi-label co-occurrence relationships. For example, Sumbul and Demir (Sumbul and Demir, 2021) combine CNN feature extraction with GNN-based similarity refinement using multiple Siamese networks and triplet loss. Wang et al. (Wang et al., 2021) propose a deep hashing approach using Graph Neural Networks (GNNs), where deep features are extracted from a pre-trained ResNet network. Other studies leverage GNNs for cross-modal matching (e.g., text-image retrieval) and multi-label image retrieval, demonstrating improvements when attention mechanisms are introduced (Yu et al., 2022; Kang et al., 2021). Contrastive learning and triplet loss have also been applied to enforce neighbourhood consistency in embedding spaces (Zhang et al., 2021; Yan et al., 2020; Draganov and Dohn, 2023). These approaches highlight the importance of preserving local context, such as adjacency to water or urban density, when organising feature spaces for retrieval and labelling. Chaudhuri et al. identified a gap between low-level features and high-level semantic concepts (Chaudhuri et al., 2018). They proposed a sub-graph matching strategy that emphasizes neighbourhood contexts within segmented images. While effective, this approach is computationally expensive (Chaudhuri et al., 2018). More recent work (Chaudhuri et al., 2022) improves on this by using contrastive loss between Siamese Graph Convolutional Networks (GCNs), replacing sub-graph matching. This method applies context-based attention to nodes, enhancing image retrieval performance. The key advantage of GCNs is their ability to encode neighbourhood information into each segment through message-passing operations, facilitating more context-aware feature extraction.
Manifold Embedding and Interactive Labelling.
Dimensionality reduction techniques such as t-SNE and UMAP have been explored to visualise high-dimensional feature spaces for interactive labelling. Halladin-Dąbrowska et al. (Halladin-Dąbrowska et al., 2019) applied t-SNE to enhance reference data for heterogeneous vegetation. Their method provides contextual feedback about the relationships between features, such as distinguishing urban areas from surrounding land or dense urban zones. This contextual discrimination, based on geographical proximity, allows labellers to customise datasets more effectively, ensuring they can balance and fine-tune their labelled datasets according to specific requirements. Patel et al. (Patel et al., 2023) proposed an interactive tool that embeds features into a 2D space for exploration and annotation. These techniques are useful for preserving local distances between data points when reducing dimensionality, which aids in visualising machine learning results (Ali et al., 2019, 2021). Their approach enables experts to label clusters visually and interactively but suffers from outliers and orientation sensitivity, where strong texture gradients (e.g., land-water boundaries) distort clustering. These limitations underscore the need for embeddings that are both context-aware and rotationally invariant.
Recent work by Sun et al. (Sun et al., 2025b) and Sun et al. (Sun et al., 2025a) introduces highly efficient frameworks for unsupervised change detection in heterogeneous and multimodal remote sensing imagery. RIEM (Sun et al., 2025b) leverages rule-based constraints within superpixels to detect changes without explicit image comparison, while LPEM(Sun et al., 2025a) preserves locality and spatial continuity through an energy-based formulation, avoiding intermediate difference images. These are highly efficient and improve robustness to imaging conditions by avoiding complex transformations for aligning heterogeneous data. Both methods share our goal of reducing reliance on predefined labels and exploiting spatial context. In contrast, our approach targets unsupervised labelling and clustering rather than change detection, learning context-aware embeddings via CNN and GNN to enable interactive exploration and semantic organisation of large datasets.
Limitations of Existing Approaches
Overall, these studies demonstrate a progression from low-level descriptors to deep learning and graph-based approaches, with increasing emphasis on contextual relationships and interactive labelling. However, existing methods still face challenges in semantic organisation, robustness to outliers, and adaptability to new classes – motivating our proposed pipeline that integrates CNNs and GNNs for context-aware, unsupervised labelling.
Existing methods face key limitations: (i) dependence on predefined label categories, reducing adaptability; (ii) loss of semantic relationships in codebook-based approaches; (iii) sensitivity to outliers and orientation, which can distort clustering; and (iv) insufficient encoding of spatial context, which is critical for remote sensing semantics. These gaps motivate the need for a more robust, context-aware solution that supports unsupervised labelling and interactive exploration.
Our Contribution
To address the challenge of costly and time-consuming labelling in remote sensing, we sought to develop a toolchain that enhances efficiency by presenting experts with visually and contextually similar content for rapid annotation. Inspired by Content-Based Image Retrieval (CBIR), which retrieves similar images through feature representation, indexing, and similarity measurement (Li et al., 2021b), we extend this concept to an unsupervised setting where semantic organisation is achieved without predefined labels. Unlike prior approaches that rely on low-level descriptors or purely deep learning-based embeddings, our method integrates Convolutional Neural Networks (CNNs) for rich feature extraction and Graph Neural Networks (GNNs) for encoding neighbourhood relationships, enabling context-aware embeddings that preserve spatial semantics. This design mitigates the impact of outliers, introduces rotational invariance, and supports interactive exploration of large datasets. In contrast to existing manifold embedding tools (Patel et al., 2023), which suffer from feature misalignment and orientation sensitivity, our pipeline provides a more robust and reliable foundation for visual labelling.
Specifically, we make the following contributions:
@@
    By integrating GNNs and feature encoding on image segmentations, we demonstrate that the resulting two-dimensional embedding space is robust to spatial transformations and encodes local context.
@@
    Finally, we apply this methodology to develop an advanced interactive tool for exploration and labelling, enhancing both efficiency and accuracy in dataset creation.
@@ 2 Materials and Methods
The proposed pipeline (Figure 1) consists of three key components: a U-Net for feature extraction, a Graph Neural Network (GNN) for generating relational embeddings, and a graph-matching algorithm for assessing graph similarity. This design explicitly bridges the gap between low-level pixel features and high-level semantic concepts as identified by Chaudhuri et al. (Chaudhuri et al., 2018). First, SLIC segmentation operates directly on raw pixel values to partition the image into super-pixels, creating spatially coherent regions that serve as graph nodes. These segments preserve local spectral and textural characteristics while introducing a structural representation of the image. Next, the U-Net extracts rich activation maps from the original image, which are aggregated within each segment to form node-level feature vectors. The GNN then propagates information across these nodes, encoding neighbourhood relationships such as adjacency to water or urban density. Through this message-passing process, isolated low-level descriptors are transformed into context-aware embeddings that capture both local detail and spatial semantics. Finally, graph matching and UMAP projection organise these enriched embeddings into a manifold where clusters correspond to meaningful categories (e.g., vegetation, urban, coastal), enabling interactive labelling at both chip and segment levels. This progression from raw pixels to graph-based context ensures that local pixel information is contextualised within its environment, achieving the bridging between low-level and high-level representations.
@@ 2.1 Dataset
We utilise the Sentinel-2 Water Edges Dataset (SWED) developed by the UK Hydrographic Office (Seale et al., 2022). It consists of globally distributed Sentinel-2 scenes captured 2017–2021 annotated for coastline extraction, although we do not use the labels. The images are tiles, regions of 10,980m by 10,980m, taken in a single pass of the satellite. Minimal cloud coverage is present within the dataset, and each tile contains a diverse set of coastal landscape features from globally distributed locations.
@@ 2.2 Fuzzy C-Means target extraction
Due to the spatial resolution of most satellite platforms, pixel assignment can be ambiguous or uncertain (Zhang et al., 2024). Fuzzy C-means extraction, unlike its K-means counterpart, assigns to each data point a degree of membership to each cluster, transforming the problem into a multi-label classification. We use all 12 spectral channels for clustering. To reduce memory requirements during model training, we apply uniform subsampling by selecting every $56^{\text{th}}$ pixel across the image, ensuring representative coverage while significantly lowering computational cost. After training, the clustering model is applied to the full-resolution image to obtain tile-wide classifications. The resultant centroids do not correspond to specific physical materials (e.g., water or concrete) but rather represent dominant spectral patterns, combinations of reflectance values across bands, that capture the most common pixel properties in the scene. These properties reflect typical spectral signatures rather than discrete land-cover classes, enabling soft classification and generalisation for pixels that do not fully belong to any single cluster.
@@ 2.3 Pre-Proccessing
For processing by the U-Net and GNN we split the tile into chips. A chip is created by applying a sliding window of size $256\times 256$ pixels to the larger image, with no overlap between consecutive windows. This results in $42\times 42$ chips containing $10,752\times 10,752$ of the original tiles $10,980\times 10,980$ pixels. For training and testing, we split the dataset by taking every fourth chip in a top-down, row-by-row pattern for the test set obtaining a 75/25 split. The size of the chip is a balance between training size and preserving contextual neighbourhood information. The larger the image the more computationally expensive the resulting pipeline becomes however more information about neighbouring features can be encoded. Considering CNN filters, larger chips result in fewer edge cases.
@@ 2.4 CNN feature extraction
To learn the texture representation we utilise a CNN where the learning function considers the C-Means classification (Bezdek et al., 1984) for an objective discrimination of features. The CNN of choice was a U-Net, with 5 layers each for the encoder and decoder. Starting with 64 kernels in the encoder, these are doubled for each consecutive layer to 1024 in the final layer and opposite for the decoder. The target of the model is the fuzzy C-Means predictions. We used 2, 8 and 18 clusters to explore the performance of the U-Net. As each fuzzy classification is non-exclusive, a multi-label loss function is used. The loss function combined Dice loss and binary cross entropy Seale et al. (2022).